A communication signal is given by the function =sin⁡t/t . Derive an equation for dy/dt using the Quotient Rule.
1
Expert's answer
2020-05-07T17:56:41-0400

Quotient Rule: "(\frac{u}{v})'=\frac{u'v-uv'}{v^2}"

"y=\frac{sin(t)}{t}"

here "u=sin(t)" and "v=t"

"u'=\frac{du}{dt}=cos(t)"

"v'=\frac{dv}{dt}=1"

So, "\frac{dy}{dt}=\frac{cos(t)\cdot t-sin(t)\cdot 1}{t^2}=\frac{t\cdot cos(t)-sin(t)}{t^2}"